OpenStudy (anonymous):

what is 5/8 divided by 11/12

Nnesha (nnesha):

\[\frac{ \frac{ 5 }{ 8 } }{ \frac{ 11 }{ 12 } }\]

Nnesha (nnesha):

\[\huge\rm \frac{ \frac{ a }{ b } }{ \frac{ c }{ d } }= \frac{ a }{ b } \times \frac{ d }{ c }\] change division to multiplication when you do that you have to flip the fraction which is at the denominator

OpenStudy (anonymous):

thanks

Nnesha (nnesha):

let me know what you get :)

OpenStudy (anonymous):

k

TheSmartOne (thesmartone):

It has been one hour, have you gotten anything so far? @jakashaka123

OpenStudy (anonymous):

srry and yes

OpenStudy (anonymous):

i had tech difficulties my laptop went out with our computer soo